3- Solve the ondulatory problem on (<2<1,t> 0 un +1:/3 = Ur. u(2.0) = 1. 2: (2.0) = 0. u,(0,t) = 0, 4,(7.t) = 0.
Answer #1


utt t ut unn OLN L32 3 19 PDE utt d = -x 2 x 80 tro ure, o)=1 ut(x,od=0 Ua(2,0-0 Unn, td=0 By method of sebarcation of variab CE TI + -२२ T x 3T X 22 on soking x(n)=c, los ant lasinda x l- & [-ci sinun & calosta] o= x lo) - ula - (2=0 0 = X(R) = a[-CE TI + -२२ T x 3T X 22 on soking x(n)=c, los ant lasinda x l- & [-ci sinun & calosta] o= x lo) - ula - (2=0 0 = X(R) = a[-) = - + - 362 - - (36) 6 ( 38 / 142 78 ) = To 0) . - 4/6 e + منو ها + ہر دل ما را در این 94ل 22 (5834 ها + ( on to J3 تي (198